OpenStudy (anonymous):

will fan & medal!!! Choose the graph that represents the equation y=x-2.

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

b.

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

d.

OpenStudy (anonymous):

Which ones do you think you can rule out?

OpenStudy (anonymous):

It's a positive slope btw

OpenStudy (anonymous):

I can rule out A and D?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

So, what do i do next?

OpenStudy (anonymous):

ok, use the equation y=x-2

OpenStudy (anonymous):

If you make x=0, what does y equal?

OpenStudy (anonymous):

it would be 2

OpenStudy (anonymous):

nope, y = (0) - 2

OpenStudy (anonymous):

isn't it still 2?

OpenStudy (anonymous):

no it's -2

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

0 - 2 =?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

ok, so now look at the graphs b and c and find the where x=0

OpenStudy (anonymous):

The graph which has a point at (0,-2) is the correct graph

OpenStudy (anonymous):

would it be c?

OpenStudy (anonymous):

That's right, good job

OpenStudy (anonymous):

thanks so much!! i fanned and gave medal! c: